OpenStudy (anonymous):

Determine the extreme values of f(x) = x + 1/ sqrt (x) . 1 < x < 9

OpenStudy (anonymous):

start with the derivative, get \(1-\frac{1}{2\sqrt{x^3}}\) set equal to zero and solve

OpenStudy (anonymous):

\[2\sqrt{x^3}-1=0\] \[\sqrt{x^3}=\frac{1}{2}\] \[x^3=\frac{1}{4}\] \[x=\frac{1}{\sqrt[3]{4}}\]

OpenStudy (anonymous):

check the critical point and the end point biggest is biggest, smallest is smallest